OpenStudy (sk12345):

An empty 330ml bottle weights 200g. If the bottle was half filled with water, would you predict that it would sink or float? use the concept of density to justify your answer. Assume that 1ml of water weights 1gram . NEEEEED HELP please

OpenStudy (irishboy123):

you need to find the density of the bottle first its mass M is \(M = \color{red}{200} + \color{green}{\frac{1}{2} \times 330 \times 1}\) grammes the red bit is its own mass. the green bit is the mass of the water inside it then its density \( \rho = \dfrac{M}{330}\) g/ml compare \(\rho\) to the density of water